OpenStudy (anonymous):

evaluate the integral from [-2,2] sqrt(4-x^2)dx

OpenStudy (anonymous):

\[2\pi\]

OpenStudy (anonymous):

method is to think that \[y=\sqrt{4-x^2}\] is the upper half of a semicircle of radius 2. area of a circle of radius 2 is \[4\pi\] and half of that is \[2\pi\]